OpenStudy (anonymous):

If f(a) = 2^a, then log base 2 f(a) = (a) 2 (b) f(a) (c) a (d) 1/2^a (e) a^2

OpenStudy (lgbasallote):

hint: \[\huge \log_x y = z \implies x^z = y\]

OpenStudy (anonymous):

so 2^x would equal 2a

OpenStudy (lgbasallote):

hmm?

OpenStudy (anonymous):

if you plug in the value you get log base 2 2a right?

OpenStudy (lgbasallote):

hmm you have \[\Large f(a) = 2^a\]correct?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

and this is in the form \[\Large y = x^z\] agree?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

so which one is a?

OpenStudy (lgbasallote):

i mean which one is x

OpenStudy (lgbasallote):

in \[\Large f(a) = 2^a\]

OpenStudy (anonymous):

oh idk i just made that the solution to the problem

OpenStudy (lgbasallote):

look closely \[\huge y = x^z\] y is the power x is the base z is the exponent so in \[\huge f(a) = 2^a\]which one is x?

OpenStudy (anonymous):

oh it's 2

OpenStudy (lgbasallote):

right. which is y?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

f(a) ?

OpenStudy (lgbasallote):

f(a) is y so which one is z?

OpenStudy (anonymous):

just a

OpenStudy (lgbasallote):

good..so x = 2 y= f(a) z = a so what is \[\huge \log_x y = z\]

OpenStudy (anonymous):

log base 2 f(a) = a

OpenStudy (lgbasallote):

right. so what's the answer to your problem?

OpenStudy (anonymous):

(c) a

OpenStudy (lgbasallote):

correct

OpenStudy (anonymous):

thank you :)

OpenStudy (lgbasallote):

welcome